# SupertonicTTS: Towards Highly Scalable and Efficient Text-to-Speech System

### v2

@@ 4.2 Optimization
We optimized the speech autoencoder for 1.5M iterations using the AdamW optimizer [36] with a learning rate of and a batch size of 128. We used four NVIDIA RTX 4090 GPUs. The loss function coefficients for training the generator were configured as follows: , , and . For adversarial training, we randomly cropped segments of real and generated speech to , following the approach in VITS [25]. The log-scaled mel spectrogram input was obtained using an FFT size of 2048 (), a Hann window of the same size, a hop size of 512 (), and 228 mel bands. The dimension of the latent space was set to 24, approximately one-tenth of the input mel spectrogram channels. Additional details on the optimization of the speech autoencoder are provided in Appendix B.1.
The text-to-latent module was optimized using the AdamW optimizer for 700k iterations with a batch size of 64 and an expansion factor . The learning rate was initially set to and halved every 300k iterations. Training was conducted on four RTX 4090 GPUs. Latents were normalized using precomputed channel-wise mean and variance statistics before being input into the text-to-latent module. The temporal compression factor was set to 6, resulting in a compressed latent dimension of 144. We set and . During training, reference speech segments were obtained by randomly cropping the input audio, with durations ranging from to . We ensured that the cropped length did not exceed half of the original speech duration.
@@ 5.1 Speech reconstruction
| test-clean | test-other | ||||||
@@
For evaluation of speech reconstruction, we compare the speech autoencoder with BigVGAN333We use the official BigVGAN-v2 checkpoint, pretrained on multilingual speech, environmental sounds, and musical instruments. [11], a state-of-the-art vocoder, using NISQA [40], UTMOSv2 [2], and CREPE [27]. Both NISQA and UTMOSv2 are non-intrusive mean opinion score (MOS) prediction systems to evaluate the perceptual quality of speech signals. Meanwhile, we use CREPE to analyze pitch in speech and compute the F1 score for voiced/unvoiced classification (V/UV F1), which serves as a key metric for identifying artifacts commonly found in non-autoregressive GAN vocoders. Furthermore, we measured the real-time factor (RTF) on an RTX 4090 GPU to quantify inference speed. Our evaluation is conducted using audio samples sourced from the test-clean and test-other subsets of LibriTTS444In this study, all evaluations were conducted using audio samples ranging from 4 to 10 seconds in length..
@@
We report the results in Table 2. In terms of NISQA scores, the speech autoencoder performs competitively, achieving 4.06 on test-clean and outperforming BigVGAN on test-other with a score of 3.76. Similarly, in terms of UTMOSv2 scores, the speech autoencoder slightly lags behind BigVGAN on test-clean but surpasses it on test-other. This indicates that the speech autoencoder can synthesize perceptually high-quality speech. For the F1 evaluation of V/UV classification, BigVGAN achieves the highest scores (0.9735 on test-clean and 0.9620 on test-other), while the speech autoencoder follows closely with 0.9587 and 0.9450, respectively. This suggests that our model introduces minor artifacts affecting voiced/unvoiced classification. Nonetheless, the overall performance remains competitive considering the V/UV F1 scores of other neural vocoders reported in the baseline paper.555The authors of BigVGAN [11] reported that HiFi-GAN [30] and WaveFlow [44] achieved V/UV F1 scores of 0.9300 and 0.9410, respectively, on the dev subsets of LibriTTS. Notably, the speech autoencoder achieves an inference speed more than 20 times faster than BigVGAN. These results collectively confirm that the speech autoencoder, despite its bottlenecked architecture and low-dimensional latent space, can generate high-quality speech while offering significantly faster inference.
5.2 Trade-off between the number of function evaluations and synthesis quality
| NFE | RTF ↓ | WER ↓ | SIM ↑ | NISQA ↑ | |
| GT | - | - | 2.181 | 0.677 ± 0.006 | 4.070 ± 0.029 |
| Ours | 4 | 0.006 | 11.43 | 0.335 ± 0.003 | 2.623 ± 0.020 |
[…]
| 128 | 0.140 | 2.693 | 0.468 ± 0.003 | 4.070 ± 0.014 |
SupertonicTTS synthesizes speech through Euler’s method during inference. By adjusting the number of function evaluations (NFE), we can balance the trade-off between synthesis speed and quality. To quantify this relationship, we generated five samples per utterance from the test-clean subset of LibriSpeech [42] by varying NFE values while keeping a CFG coefficient to 3. These samples were then evaluated in terms of RTF, word error rate (WER), NISQA score [40], and speaker similarity (SIM). SIM was calculated as the cosine similarity between speaker embeddings from the generated speech and the corresponding 3-second reference, using the WavLM-TDNN model [7].
The evaluation results, presented in Table 3, demonstrate a general trend that increasing NFE leads to improvements in the WER, SIM, and NISQA scores. Specifically, the best scores for WER and SIM were obtained at NFE values of 32 and 16, respectively. This suggests that NFE values exceeding 32 effectively capture intelligibility, prosodic naturalness, and speaker identity. Meanwhile, NISQA scores exhibited a consistent positive correlation with NFE, suggesting that higher NFE values result in enhanced audio fidelity. However, this improvement comes at the cost of increased processing time, as reflected in the RTF. Given this trade-off, we proceed with an NFE of 32 for further evaluations.
@@ 5.3 Evaluation of context-sharing batch expansion
To assess the impact of the context-sharing batch expansion on training efficiency, we trained four additional SupertonicTTS models using the following batch size () and expansion factor () combinations: (64, 1), (128, 1), (256, 1), and (64, 2). All other architectural and optimization settings remained unchanged. Performance was assessed using two metrics: validation loss and automatic speech recognition (ASR) results. Validation loss was computed on the test-clean subset of LibriTTS by averaging flow-matching losses (Eq. 1) across five timesteps: 0.1, 0.3, 0.5, 0.7, and 0.9. To ensure consistency, the reference mask was kept identical across all evaluations. For evaluation on ASR results, we generated five samples per utterance from the test-clean subset of LibriSpeech, using a 3-second reference speech segment cropped from a different utterance of the same speaker. To address training–inference mismatch due to the absence of punctuation in LibriSpeech, we utilized LibriSpeech-PC [39] for model input. If a corresponding transcription was not available in LibriSpeech-PC, the original LibriSpeech text was capitalized and appended with a period. The generated samples were transcribed using the CTC-based HuBERT-Large model666https://huggingface.co/facebook/hubert-large-ls960-ft. [16] and WER was calculated by comparing the ASR outputs to the original LibriSpeech transcriptions. All transcriptions were normalized using NVIDIA’s NeMo-text-processing [63, 4] before measuring WER.
We present validation loss and WER curves throughout training up to 300k iterations in Figure 2. It can be observed that the validation loss converges faster as the expansion factor increases. This trend is very similar to the effect of increasing the batch size . Interestingly, we found that WER improves more quickly with an increase in the expansion factor, while increasing the batch size leads to degraded WER performance. We hypothesize that this is because the proposed method alleviates the speech-text alignment issue by using the same text and speech pair with different noise and time samples, and simply increasing the number of speech-text pairs does not effectively improve alignment learning. Evaluation results throughout the entire training process are provided in Appendix D.1.
We also assessed the computational efficiency of context-sharing batch expansion by analyzing memory usage, time per iteration, and the number of floating-point operations (GFLOPs). For testing, we used 15-second input speech, 250 characters, and 3-second reference speech. Iteration times were measured over 100 trials using a single RTX 4090 GPU, with 95% confidence intervals narrower than . Table 4 presents the results for varying batch sizes and expansion factors . We observed that increasing consistently offered better efficiency compared to increasing across all metrics. Notably, raising from 1 to 4 resulted in a 64% increase in iteration time, whereas increasing by a factor of 4 led to a 253% increase . The overall results demonstrate that the proposed method improves training efficiency in memory usage and processing time by sharing conditional variables across batch data expanded from the same source.
@@ 5.4 Comparison with other zero-shot TTS models
| WER ↓ | CER ↓ | SIM ↑ | Inf. Time ↓ | Data (hour) | #Param. | |
| GT | 2.18 | 0.60 | 0.6770 | - | - | - |
| VALL-E | 5.9 | - | - | 6.4s | 60,000 | 410M |
| VoiceBox | 1.9 | - | 0.662 | 6.2s | 60,000 | 371M |
[…]
| CLaM-TTS | 5.11 | 2.87 | 0.495 | 4.15s (A100) | 55,000 | >1.3B |
[…]
| Ours | 2.64 | 0.83 | 0.472 / 0.915† | 0.23s (RTX 4090) | 945 | 44M |
In this section, we benchmark our proposed model with state-of-the-art TTS systems. Specifically, we evaluate our model against five strong baselines: VALL-E [55], VoiceBox [32], Mega-TTS 2 [21], CLaM-TTS [26], and DiTTo-TTS [33]. These models exhibit impressive TTS performance, even in zero-shot scenarios. For the evaluation of SupertonicTTS, we followed the same experimental setup as described in Section 5.3. Additionally, we computed character error rates (CER) to assess the finer details of pronunciation in the synthesized speech. To compare speaker similarity with Mega-TTS 2, we also employed WavLM Base+777https://huggingface.co/microsoft/wavlm-base-plus-sv. used in the Mega-TTS 2 paper. Inference time was measured by averaging the processing time for 100 generations of 10-second speech, performed on an RTX 4090.
Table 5 presents the evaluation results, with baseline scores sourced from their publications. In terms of WER, SupertonicTTS performed competitively with a score of 2.64, surpassing VALL-E (5.9) and CLaM-TTS (5.11). For CER, SupertonicTTS achieved a strong result of 0.83, outperforming DiTTo-TTS (0.89). These results suggest that SupertonicTTS can generate speech with minimal skipping, repetition, or mispronunciation issues. Regarding speaker similarity, as measured by WavLM-TDNN, SupertonicTTS scored 0.472 which is lower than VoiceBox (0.662), CLaM-TTS (0.495), and DiTTo-TTS (0.627). This difference may stem from SupertonicTTS’s lack of use of text information for reference speech unlike the other models, potentially affecting zero-shot performance. However, SupertonicTTS achieved a score of 0.915 when evaluated with WavLM Base+, surpassing Mega-TTS 2, which also does not use text for reference speech. Notably, these results were achieved with significantly fewer parameters and less than 2% of TTS dataset size compared to the baselines. Lastly, SupertonicTTS exhibited exceptional inference speed, requiring just 0.23 seconds to generate a 10-second speech sample. This result is considerably faster than the next fastest model, DiTTo-TTS, which requires 1.62 seconds. Overall, SupertonicTTS demonstrates competitive performance with remarkable efficiency in both architecture and training. A subjective listening test in Appendix D.2 further confirms its perceptual competitiveness in speech naturalness and speaker similarity.
@@ 6 Conclusion
In this paper, we introduced SupertonicTTS, a novel text-to-speech system designed to effectively address the limitations of contemporary TTS models. By leveraging the LDM framework, we designed a system comprising a speech autoencoder, flow-matching-based text-to-latent module, and utterance-level duration predictor. To enhance scalability and efficiency, we incorporated a low-dimensional latent space, latent compression, context-sharing batch expansion, and ConvNeXt blocks. Furthermore, we simplified the pipeline by eliminating external dependencies such as G2P modules, text-speech aligners, and pretrained text encoders. Our extensive experiments validated that SupertonicTTS provides competitive zero-shot TTS performance with only 44 million parameters and fast inference speed. We believe the proposed system substantially reduces computational complexity and overhead in speech synthesis, opening promising avenues for future research in diverse speech applications and real-time scenarios.
@@ A.1 Speech autoencoder
We illustrate the detailed architecture of the speech autoencoder in Figure 3. Both the latent encoder and the latent decoder are built on the Vocos architecture [51] with several modifications, aiming for useful applications (e.g., latent encoding, fast inference, and low-latency TTS).
@@ A.2 Text-to-latent module
We illustrate the detailed architecture of the text-to-latent module in Figure 4. Each component is carefully designed to operate efficiently with a simple architecture. Note that the text-to-latent module does not rely on any external pretrained models, G2P modules, or text-speech aligners.

### v3

@@ 4.2 Optimization
We optimized the speech autoencoder for 1.5M iterations using the AdamW optimizer [42] with a learning rate of and a batch size of 128. We used four NVIDIA RTX 4090 GPUs. The loss function coefficients were configured as follows: , , and . For adversarial training, we randomly cropped segments of real and generated speech to . The log-scaled mel spectrogram input was obtained using an FFT size of 2048 (), a Hann window of the same size, a hop size of 512 (), and 228 mel bands. Additional details on the optimization of the speech autoencoder are provided in Appendix B.1.
The text-to-latent module was optimized using the AdamW optimizer for 700k iterations with a batch size of 64 and an expansion factor . The learning rate was initially set to and halved every 300k iterations. Training was conducted on four RTX 4090 GPUs. Latents were normalized using precomputed channel-wise mean and variance statistics before being input into the text-to-latent module. We set and . During training, reference speech segments were obtained by randomly cropping the input audio, with durations ranging from to . We ensured that the cropped length did not exceed half of the original speech duration.
@@ 5 Experiments
We used four test sets for evaluation, each containing audio samples ranging from 4 to 10 seconds: LT-clean, LT-other, LS-clean, and LS-PC-clean. LT-clean and LT-other were derived from the test-clean and test-other sets of LibriTTS [73], respectively. LS-clean was sourced from the test-clean set of LibriSpeech [49], while LS-PC-clean corresponds to the test set proposed by Chen et al. [7]. For the text-to-latent module, we set the number of function evaluations (NFE) to 32, with its effect analyzed in Appendix D.1.
@@ 5.1 Speech reconstruction
| LT-clean | LT-other | ||||||
@@
We evaluated reconstruction quality by comparing the speech autoencoder with the official checkpoint of BigVGAN [38]. Experiments were conducted on LT-clean and LT-other using three metrics: NISQA [45], UTMOSv2 [2], and CREPE [31]. Both NISQA and UTMOSv2 are mean opinion score (MOS) prediction systems to estimate the perceptual quality of speech signals. CREPE is used to analyze pitch in speech and compute the F1 score for voiced/unvoiced classification (V/UV F1), which serves as a key metric for identifying artifacts in generated speech. Additionally, we measured the real-time factor (RTF) on an RTX 4090 GPU to quantify inference speed.
@@
We report the results in Table 2. In terms of NISQA scores, the speech autoencoder performs competitively, achieving 4.06 on LT-clean and outperforming BigVGAN on LT-other with a score of 3.76. Similarly, in terms of UTMOSv2 scores, the speech autoencoder slightly lags behind BigVGAN LT-clean but surpasses it on LT-other. These results support that the speech autoencoder can synthesize perceptually high-quality speech. For the F1 evaluation of V/UV classification, BigVGAN achieves the highest scores (0.9735 on LT-clean and 0.9620 on LT-other), while the speech autoencoder follows closely with 0.9587 and 0.9450, respectively. This suggests that our model introduces minor artifacts affecting voiced/unvoiced classification. Nonetheless, the overall performance remains competitive considering the V/UV F1 scores of other neural vocoders reported in the baseline paper.111Lee et al. [38] state that HiFi-GAN [35] and WaveFlow [51] achieve V/UV F1 scores of 0.9300 and 0.9410, respectively, on the dev subsets of LibriTTS. Notably, the speech autoencoder achieves an inference speed more than 20 times faster than BigVGAN. These results collectively confirm that our speech autoencoder, despite its bottlenecked architecture and low-dimensional latent space, can generate high-quality speech while offering a substantial advantage in inference efficiency.
@@ 5.2 Evaluation of context-sharing batch expansion
| WER(%) | CER(%) | ||
| 64 | 1 | 3.11 | 1.08 |
| 128 | 1 | 3.08 | 1.06 |
| 64 | 2 | 2.97 | 1.00 |
| 256 | 1 | 2.88 | 0.92 |
| 64 | 4 | 2.64 | 0.83 |
To assess the effectiveness of context-sharing batch expansion, we trained four additional models using the following combinations of batch size and expansion factor : (64, 1), (128, 1), (256, 1), and (64, 2). Performance was evaluated with validation loss and pronunciation error. Validation loss was computed on LT-clean by averaging across five timesteps: 0.1, 0.3, 0.5, 0.7, and 0.9. Pronunciation error was quantified using word error rate (WER) and character error rate (CER) between synthesized speech transcriptions and the ground-truth. For this evaluation, we generated five samples per utterance from LS-clean and transcribed them using a CTC-based HuBERT-Large model [20]. All transcriptions were normalized using NVIDIA’s NeMo-text-processing [74] before computing WER and CER.
@@
Fig. 2 shows validation loss and WER curves throughout training up to 300k iterations. It can be observed that the validation loss converges faster as increases, similar to the effect of increasing . Interestingly, WER improves more rapidly with higher , whereas increasing actually degrades performance, particularly in the early stages of training (e.g., around 100k iterations). Although this gap narrows as training progresses, models trained with larger ultimately achieve lower final WER and CER than those with correspondingly larger , as summarized in Table 3. We hypothesize that presenting the same text-speech pair with varying noise and timesteps is more effective for alignment learning compared to using different text–speech pairs. Complete evaluation curves for the entire training process are provided in Appendix D.2.
We also assessed the computational efficiency of context-sharing batch expansion by analyzing GPU memory usage (GiB), time per iteration (seconds), and the number of floating-point operations (GFLOPs). For testing, we used 15-second input speech, 250 characters, and 3-second reference speech. Iteration times were measured over 100 trials using a single RTX 4090 GPU, with 95% confidence intervals narrower than . GPU memory usage and iteration time were measured for a single training iteration, while GFLOPs were calculated from a single forward pass through the text-to-latent module. Table 4 shows that increasing consistently offers better efficiency compared to increasing across all metrics. Notably, raising from 1 to 4 results in a 64% increase in iteration time, whereas increasing by a factor of 4 leads to a 253% increase. The overall results demonstrate that the proposed method not only stabilize text-speech alignment but also significantly enhances training efficiency with respect to memory usage and processing time.
@@ 5.3 Comparison with other zero-shot TTS models
| Test set | Model | WER | CER | Data | #DP | #T2F | #F2S | #All | RTF |
| LS-clean | GT | 2.18 | 0.60 | - | - | - | - | - | - |
| VALL-E | 5.9 | - | 60k | - | 403M† | 7M | 410M† | 0.64 | |
| VoiceBox | 1.9 | - | 60k | 28M | 330M | 13M | 371M | 0.62 | |
| CLaM-TTS | 5.11 | 2.87 | 55k | - | >1.23B† | 112M | >1.3B† | 0.42 (A100) | |
| DiTTo-TTS | 2.56 | 0.89 | 55k | 33M | 825M | 112M | 940M | 0.16 (A100) | |
| Ours | 2.64 | 0.83 | 945 | 0.5M | 18.5M | 25M | 44M | 0.02 (RTX 4090) | |
| LS-PC-clean | GT | 1.86 | 0.50 | - | - | - | - | - | - |
| FireRedTTS | 2.69 | - | 248k∗ | - | 538M | 235M | 773M | 0.84 (RTX 3090) | |
| F5-TTS | 2.42 | - | 100k∗ | - | 335.8M | 13.5M | 349M | 0.31 (RTX 3090) | |
| Ours | 2.41 | 0.80 | 945 | 0.5M | 18.5M | 25M | 44M | 0.05 (RTX 3090) |
In this section, we provide a comparative analysis of our proposed model against state-of-the-art TTS systems. Specifically, we evaluate our model against six baselines: VALL-E [67], VoiceBox [36], CLaM-TTS [30], DiTTo-TTS [37], FireRedTTS [14], and F5-TTS [7]. These models exhibit exceptional TTS performance even in zero-shot scenarios, establishing them as strong benchmarks for our study.
Table 5 presents an overall comparison of pronunciation errors (WER and CER), the amount of transcribed speech used for training (Data), model parameter counts (#DP, #T2F, #F2S, and #All), and inference speed (RTF). Baseline results are sourced from their publications. On the LS-clean benchmark, SupertonicTTS achieves a WER of 2.64 and the lowest CER of 0.83, demonstrating strong pronunciation accuracy. The performance is competitive with much larger systems such as DiTTo-TTS (WER 2.56, CER 0.89) and VoiceBox (WER 1.9), despite of much smaller parameter count. Also, our model achieves the lowest WER of 2.41 on LS-PC-clean, outperforming FireRedTTS (WER 2.69) and F5-TTS (WER 2.42). While data efficiency is not the primary focus of this study, it is noteworthy that our system achieves such results with a training corpus that is orders of magnitude smaller than those used by other systems. From the perspective of model size, SupertonicTTS is remarkably compact. The entire system contains only 44M parameters, compared to hundreds of millions or even billions in the baselines. In particular, our text-to-latent module (#T2F) requires 18.5M parameters, whereas the next smaller model, VoiceBox, allocates 330M parameters, which is approximately 18 times larger. This parameter efficiency highlights the architectural effectiveness of our design in capturing linguistic and acoustic mappings with minimal overhead. Another key strength of our system lies in inference speed. With an RTF of 0.02 on an RTX 4090 and 0.05 on an RTX 3090, our model runs substantially faster than all baselines, which report RTFs between 0.16 and 0.84. Such efficiency makes our approach well-suited for real-time or resource-constrained deployment scenarios, where large-scale TTS systems often face limitations.
| Preference test | Reference quality | |||
| Naturalness | Similarity | PESQ | SI-SDR | |
| Ours vs. VALL-E | 0.233 ± 0.112 | 0.043 ± 0.115 | 3.422 ± 0.327 | 22.526 ± 2.884 |
| Ours vs. VoiceBox | -0.476 ± 0.106 | 0.316 ± 0.113 | 2.171 ± 0.377 | 10.978 ± 8.851 |
| Ours vs. CLaM-TTS | 0.084 ± 0.106 | 0.076 ± 0.113 | 3.642 ± 0.359 | 23.071 ± 3.789 |
| Ours vs. DiTTo-TTS | 0.076 ± 0.109 | 0.057 ± 0.112 | 3.746 ± 0.253 | 24.799 ± 2.034 |
We also conducted subjective listening tests via Amazon Mechanical Turk to assess perceptual quality. Specifically, we compared SupertonicTTS against VALL-E, VoiceBox, CLaM-TTS, and DiTTo-TTS using audio samples from their respective demo pages. For each comparison, we performed preference tests on 15 paired samples with 42 participants. Each pair consisted of one sample from a baseline model and one from SupertonicTTS. Participants evaluated the pairs based on two criteria: naturalness (which sample sounds more natural) and speaker similarity (which sample sounds more similar to the reference speech). For quantitative analysis, responses were recorded on a five-point scale from -2 to +2, where positive scores indicate a preference for SupertonicTTS. Detailed instructions provided to the participants are included in Appendix E.
Table 6 summarizes the results and shows that SupertonicTTS delivers highly competitive performance. For naturalness, SupertonicTTS was clearly preferred over VALL-E (0.233) and slightly preferred over CLaM-TTS (0.084) and DiTTo-TTS (0.076). However, the subjects perceived VoiceBox samples as more natural than those from SupertonicTTS (-0.476). In terms of speaker similarity, SupertonicTTS was strongly favored over VoiceBox (0.316) and slightly outperformed the other baselines. The contrasting results relative to VoiceBox prompted further investigation. We analyzed the reference speech samples used in the subjective listening test. Specifically, we measured perceptual evaluation of speech quality (PESQ) [57] and scale-invariant signal-to-distortion ratio (SI-SDR) of these samples, which is also reported in Table 6. The analysis revealed that VoiceBox’s reference audio had significantly lower PESQ (2.171) and SI-SDR (10.978) than the comparable-quality references for the other baselines (PESQ 3.4 and SI-SDR 22). Based on these findings, we conjecture that SupertonicTTS reproduces the characteristics of a given reference more faithfully than VoiceBox, but the low reference quality likely caused the SupertonicTTS output to be judged as less natural, despite its strong speaker similarity. Overall, these experimental results support that SupertonicTTS achieves competitive performance on par with recent state-of-the-art zero-shot TTS models. We encourage readers to visit our demo page222https://supertonictts.github.io/. for a perceptual comparison.
@@ 6 Conclusion
In this paper, we introduced SupertonicTTS, a novel text-to-speech system designed to effectively address the limitations of contemporary TTS models. Within the LDM framework, we designed a system comprising a speech autoencoder, flow-matching-based text-to-latent module, and utterance-level duration predictor. To enhance efficiency and simplicity, we incorporated a low-dimensional latent space, latent compression, context-sharing batch expansion, and ConvNeXt blocks. Furthermore, we simplified the pipeline by eliminating external dependencies such as G2P modules, text-speech aligners, and pretrained text encoders. Our extensive experiments validated that SupertonicTTS provides competitive zero-shot TTS performance with only 44 million parameters and fast inference speed. We believe the proposed system substantially reduces architectural complexity and computational overhead in speech synthesis, opening promising avenues for future research in diverse speech applications and real-time scenarios.
@@ A.1 Speech autoencoder
We illustrate the detailed architecture of the speech autoencoder in Fig. 3. Both the latent encoder and the latent decoder are built on the Vocos architecture [61] with several modifications, aiming for useful applications (e.g., latent encoding, fast inference, and low-latency TTS).
@@ A.2 Text-to-latent module
We illustrate the detailed architecture of the text-to-latent module in Fig. 4. Each component is carefully designed to operate efficiently with a simple architecture. Note that the text-to-latent module does not rely on any external pretrained models, G2P modules, or text-speech aligners.